# Generalisation in Multi-task Offline Q-Learning

### v1

@@ 4.2 Theoretical Results
To facilitate the analysis, we adopt a realizability assumption that is standard in the low-rank Q-learning literature.prog
Assumption 3.1 (Low-rank realizability). There exists a shared representation $\phi^{*}\in\Phi$ and task- and stage-specific weight vectors $w_{h}^{(*,t)}\in\Psi$ such that, for all $h\in[H]$ and $t\in[T]$,
@@
For simplicity assume that $\Phi$ and $\Psi$ are finite sets. The main theorem here is the following:
@@
Under the low-rank MDP structure (Assumption 3.1), we establish finite-sample generalization bounds for multitask offline Q-learning.
@@
    This bound exhibits the expected $O(1/\sqrt{nT})$ statistical scaling due to pooling data across tasks.
@@
5 Generalization Bounds for Infinite Hypothesis Sets
When the hypothesis spaces for the encoder $\Phi$ and the decoders $\Psi$ are infinite, such as those parameterized by continuous function approximators, the Union Bound approach utilized in Theorem 1 is no longer applicable. We must instead rely on uniform convergence theory based on the Rademacher complexity to quantify the capacity of the Multitask Hypothesis Space $\mathcal{F}=\Phi\times\Psi^{T}$.
5.1 Multitask Loss Class and Rademacher Complexity
@@
$\mathcal{R}(\mathcal{G})=\mathbb{E}_{\mathcal{D},\mathbf{\sigma}}\left[\sup_{g\in\mathcal{G}}\frac{1}{nT}\sum_{t=1}^{T}\sum_{i=1}^{n}\sigma_{i}^{(t)}g(s_{h}^{(i,t)},a_{h}^{(i,t)})\right]$
@@
5.2 Generalization Bound via Rademacher Complexity
By applying the standard generalization theorem (e.g., using McDiarmid’s inequality and the Rademacher complexity bound), we can replace the cardinality-dependent term in Theorem 3.1 with the complexity measure $\mathcal{R}(\mathcal{G})$.
@@ 6 Experiments
Figure 1: T-Scaling: Proof of Multi-task Efficiency. Bound: $O(1/\sqrt{nT})$. As the number of tasks increases (with fixed $n=200$, $H=5$), the average MSE to the true $Q^{*}$ decreases consistently, confirming the multitask efficiency predicted by our bounds.
Figure 2: n-Scaling: Proof of Sample Consistency. Bound: $O(1/\sqrt{n})$. With fixed $T=5$ and $H=5$, increasing samples per task $n$ leads to a clear downward trend in error, consistent with our statistical rates despite natural training variance.
Figure 3: H-Scaling: Proof of Recursive Error Propagation. Bound: $O(H^{3}C_{\max})$. For fixed $T=5$ and $n=500$, the total statistical error grows smoothly and polynomially with horizon $H$, validating the sharpness of our Bernstein-based recursive analysis over looser alternatives.
When varying the number of tasks $T$ while fixing the number of samples per task $n$ and the horizon $H$, we observe a clear decrease in estimation error as $T$ increases. Since the total amount of offline data grows linearly with $T$ (total samples =nT=nT), Theorem 1(b) predicts a statistical term scaling as $O(1/\sqrt{nT})$, which reduces to $O(1/\sqrt{T})$ in this regime. After unrolling the recursion in Theorem 1(c), the dominant contribution to the suboptimality gap $\Delta_{1}$ scales as $O(H^{2}\lambda_{\max}/\sqrt{nT})$. The observed downward trend in error as $T$ grows is consistent with this predicted behavior and illustrates the benefit of pooling data across tasks when learning a shared representation.
@@
Finally, when varying the horizon $H$ while holding $T$ and $n$ fixed, we observe polynomial growth in the estimation error. This behavior is consistent with the horizon-dependent terms in Theorem 1(c), and rules out exponential error propagation. Overall, these experiments illustrate that multitask offline Q-learning exhibits the qualitative scaling behavior predicted by the theory with respect to the number of tasks, samples per task, and planning horizon.
@@ 7 Conclusion
We studied multitask offline reinforcement learning in settings where tasks share a common low-rank representation of their action-value functions. Focusing on a model-free, value-based approach, we analyzed a multitask variant of fitted Q-iteration that jointly learns a shared representation and task-specific decoders from offline data. Under standard realizability and coverage assumptions, we derived finite-sample generalization guarantees that explicitly characterize how pooling data across tasks improves statistical efficiency.
A key takeaway from our analysis is that multitask offline Q-learning achieves final suboptimality bounds that are competitive with existing theoretical results for both offline and multitask reinforcement learning. In particular, despite operating in a strictly more challenging setting—combining offline learning, multitask representation sharing, and model-free Bellman regression—our bounds match the best-known polynomial dependence on the horizon and exhibit the expected $1/\sqrt{nT}$ scaling due to multitask data aggregation. This suggests that multitask structure can be exploited effectively without incurring additional statistical penalties relative to single-task or model-based approaches.
We further analyzed a downstream offline learning setting in which a novel task shares the representation learned from upstream multitask data. Our results show that reusing this representation reduces the effective complexity of downstream learning and leads to improved sample efficiency compared to learning from scratch, while preserving the same order of horizon dependence.
Taken together, our findings indicate that multitask offline representation learning via value-based methods is not only theoretically sound, but also statistically competitive with existing approaches. Rather than improving rates through stronger assumptions or algorithmic pessimism, our analysis demonstrates that carefully exploiting shared structure suffices to recover the same order-optimal guarantees in a model-free framework. We view this work as a step toward a unified understanding of multitask and offline reinforcement learning, and an invitation to further investigate lower bounds, alternative coverage conditions, and extensions to richer function approximation classes.
@@
MORL ([11] ) Offline, multitask $\tilde{O}\!\left({H^{2}\lambda}/{\sqrt{nT}}\right)$
REFUEL ([5]) Offline / online, multitask $\tilde{O}\!\left(\varepsilon_{\text{rep}}+{H^{2}\lambda}/{\sqrt{n}}\right)$
Pessimistic Q-learning ([12]) Offline, single-task $\tilde{O}\!\left({H^{2}\lambda}/{\sqrt{n}}\right)$
@@ Appendix A Proof of Theorem 1a
where $\hat{\phi},\hat{P}^{(1)},\ldots,\hat{P}^{(T)}$ are the output of Algorithm 1 and an appropriete constant $B$.
@@ Appendix B Proof of Theorem 1b
We would now look at the individual terms, in the RHS:
[Explain $\lambda_{\max}$ insertion please]
@@ Appendix C Proof of Theorem 1c
Appendix D New Proof of theorem-2
@@ Appendix E Note on Raedemacher Complexity
To facilitate a deeper understanding of the bounds in Theorem 2, which leverage Rademacher complexity for sharper statistical rates in the downstream task, we provide a brief overview of Rademacher complexity. This measure of function class complexity is a key tool in statistical learning theory, offering tighter generalization bounds than alternatives like VC dimension, especially for non-i.i.d. or structured data settings common in reinforcement learning.
@@ E.2 Generalization Bounds Using Rademacher Complexity
A fundamental result (e.g., from Bartlett and Mendelson, 2002) states that, with high probability $1-\delta$, for any $g\in\mathcal{G}$ and bounded loss $\ell:[0,1]$,
@@
In summary, Rademacher complexity enables variance-aware, instance-dependent bounds, making it ideal for our sharpened downstream analysis where upstream multitasking effectively lowers the class complexity for the new task. For further reading, see [18] or [22].
@@ Appendix F Role of $\lambda_{\max}$ in the Proof
The coefficient $\lambda_{\max}$ (or $\lambda_{\max,\sup}$) appears in the error propagation step of the proof (Theorems 1(b)–1(c) and similarly in Theorem 2) to account for **distribution shift** between the offline data distribution $\mu_{b}$ (induced by the behavior policy that collected the datasets) and the distributions naturally arising under the Bellman operator (i.e., the next-state occupancies induced by applying the optimal Bellman backup).
@@
In an ideal on-policy setting, the Bellman optimality operator $T_{h}^{*}$ is a $\gamma$-contraction in the $L_{2}$ norm under the same measure in which we evaluate the error. However, in offline RL, we only have access to data from a fixed behavior policy, so the state-action pairs $(s,a)$ are distributed according to $\mu_{b}$, while the Bellman target involves expectations over next states $s^{\prime}\sim P_{h}(\cdot|s,a)$. This creates a **change of measure**: the propagated error $\|\hat{Q}_{h+1}-Q_{h+1}^{*}\|_{L_{2}(\mu_{b})}^{2}$ can be larger than the “natural” contraction would suggest.
Formally, when bounding
$\|T_{h}^{*}\hat{Q}_{h+1}-T_{h}^{*}Q_{h+1}^{*}\|_{L_{2}(\mu_{b})}^{2}\leq\|\hat{Q}_{h+1}-Q_{h+1}^{*}\|_{L_{2}(\mu_{b})}^{2}$
after applying the raw contraction, we must adjust for the fact that the $L_{2}(\mu_{b})$ norm weights states differently from the on-policy/next-step distribution. The concentrability coefficient $\lambda_{\max}$ (typically defined as the supremum over $h$ of the largest eigenvalue of the Radon–Nikodym derivative or a bounded density ratio $\sup\|dP_{h}(\cdot|s,a)/d\mu_{b}\|_{\infty}$) provides exactly this adjustment:
$\|\hat{Q}_{h+1}-Q_{h+1}^{*}\|_{L_{2}(\mu_{b})}^{2}\leq\lambda_{\max}\|\hat{Q}_{h+1}-Q_{h+1}^{*}\|_{L_{2}(\text{on-policy})}^{2},$
ensuring the propagated error does not exceed $\lambda_{\max}$ times the contracted term.
Without $\lambda_{\max}$, the bound would only hold under perfect coverage ($\mu_{b}$ visiting all relevant state-action pairs induced by optimal policies), which is unrealistic in offline settings. Including $\lambda_{\max}$ makes the analysis robust to arbitrary distribution shift, at the cost of a multiplicative factor that reflects the worst-case misalignment between offline data and the regions explored by the learned policy. This is standard in offline RL theory (e.g., [12]) and ensures our guarantees remain valid under the single-policy concentrability assumption common in the literature. In practice, when the behavior policy provides reasonable coverage, $\lambda_{\max}$ is bounded, preserving polynomial scaling in horizon and sample complexity.

### v2

@@ 4.2 Theoretical Results
The notations, ideas in this section are inspired by Borsa et al. (2016); Ishfaq et al. (2024). To facilitate the analysis, we adopt a realizability assumption Ishfaq et al. (2024); Agarwal et al. (2020a) that is standard in the low-rank Q-learning literature. Similar variations of the low-rank assumptions can be found in the literature Agarwal et al. (2020a); Oprescu et al. (2024).
Assumption 3.1 (Low-rank realizability) Agarwal et al. (2020a); Ishfaq et al. (2024). There exists a shared representation $\phi^{*}\in\Phi$ and task- and stage-specific weight vectors $w_{h}^{(*,t)}\in\Psi$ such that, for all $h\in[H]$ and $t\in[T]$,
@@
For simplicity, assume that $\Phi$ and $\Psi$ are finite sets. The main theorem here is the following:
@@
Under the low-rank MDP structure (Assumption 3.1), we establish finite-sample generalization bounds for multitask offline Q-learning Borsa et al. (2016).
@@
4.3 Generalization Bounds for Infinite Hypothesis Sets
When the hypothesis spaces for the encoder $\Phi$ and the decoders $\Psi$ are infinite, such as those parameterized by continuous function approximators, the Union Bound approach utilized in Theorem 1 is no longer applicable. We must instead rely on uniform convergence theory based on the Rademacher complexity Bartlett and Mendelson (2002); Shalev-Shwartz and Ben-David (2014); Koltchinskii and Panchenko (2002) to quantify the capacity of the Multitask Hypothesis Space $\mathcal{F}=\Phi\times\Psi^{T}$.
4.3.1 Multitask Loss Class and Rademacher Complexity
@@
$\mathcal{R}(\mathcal{G})=\mathbb{E}_{\mathcal{D},\mathbf{\sigma}}\left[\sup_{g\in\mathcal{G}}\frac{1}{nT}\sum_{t=1}^{T}\sum_{i=1}^{n}\sigma_{i}^{(t)}g(s_{h}^{(i,t)},a_{h}^{(i,t)},y_{h}^{(i,t)})\right]$
@@
4.3.2 Generalization Bound via Rademacher Complexity
By applying the standard generalization theorem (e.g., using McDiarmid’s inequality Mohri et al. (2018a), and the Rademacher complexity bound Bartlett and Mendelson (2002); Mohri et al. (2018b)), we can replace the cardinality-dependent term in Theorem 1 (c) with the complexity measure $\mathcal{R}(\mathcal{G})$.
@@ 4.3.2 Generalization Bound via Rademacher Complexity
4.4 Bounds under Non-Realisability scenarios
Our theoretical analysis assumes realizability of the underlying MDP within the chosen function class. When this assumption 4.2 is violated, our guarantees for 1 extend in a standard manner by incurring an additional approximation error term. In particular, if the true transition dynamics admit an $\epsilon$- approximation within the model class, then the resulting policy suboptimality increases by at most $\mathcal{O}(\mathrm{poly}(H)\epsilon_{\text{Approx}})$ Munos and Szepesvári (2008); Antos et al. (2008). Empirically, we observe that the proposed method remains stable under moderate misspecification, suggesting robustness beyond the realizable regime.
@@ 5 Experiments
Figure 1: Scaling behavior of the proposed method. Left: Net Multitask efficiency $(nT\times MSE)$ with increasing $T$ (fixed $n=200,H=5$). Middle: Sample consistency with increasing $n$ (fixed $T=5,H=5$). Right: Polynomial horizon dependence with increasing $H$ (fixed $T=5,n=500$).
When varying the number of tasks $T$ while fixing the number of samples per task $n$ and the horizon $H$, we observe a clear decrease in estimation error as $T$ increases. Since the total amount of offline data grows linearly with $T$ (total samples =nT=nT), Theorem 1(b) 1 predicts a statistical term scaling as $O(1/\sqrt{nT})$, which reduces to $O(1/\sqrt{T})$ in this regime. After unrolling the recursion in Theorem 1(c), the dominant contribution to the suboptimality gap $\Delta_{1}$ scales as $O(H^{2}\lambda_{\max}/\sqrt{nT})$. The observed downward trend in error as $T$ grows is consistent with this predicted behavior and illustrates the benefit of pooling data across tasks when learning a shared representation.
@@
Finally, when varying the horizon $H$ while holding $T$ and $n$ fixed, we observe polynomial growth in the estimation error. This behavior is consistent with the horizon-dependent terms in Theorem 1(c) 1, and rules out exponential error propagation. Overall, these experiments illustrate that multitask offline Q-learning exhibits the qualitative scaling behavior predicted by the theory with respect to the number of tasks, samples per task, and planning horizon.
@@ 6 Conclusion
In this work, we examined multitask offline reinforcement learning in settings where tasks share a common low-rank structure in their action-value functions. We focused on a model-free, value-based approach based on Multitask Fitted Q-Iteration (MFQI) Borsa et al. (2016), and described how it can be formulated to jointly learn a shared representation along with task-specific value functions from offline data. Under standard realizability and coverage assumptions, we discussed the corresponding finite-sample guarantees and how they capture the effect of pooling data across tasks. In particular, these results illustrate how shared representations can improve statistical efficiency in the multitask offline setting. Our presentation places MFQI in the context of existing theoretical work and highlights its connection to prior analyses of multitask representation learning in reinforcement learning Cheng et al. (2022b); Ishfaq et al. (2024); Agarwal et al. (2020a).
A key takeaway from our analysis is that the proposed MFQI-based approach attains comparable suboptimality bounds, in order, to existing theoretical results for multitask offline reinforcement learning Agarwal et al. (2020a); Ishfaq et al. (2024) under related structural assumptions 4.2 Oprescu et al. (2024); Ishfaq et al. (2024); Domingues et al. (2021). In particular, under the low-rank realizability 4.2 and coverage conditions considered here, our bounds exhibit polynomial dependence on the horizon and scale as $O(1/\sqrt{nT})$ 1 with the total number of samples, reflecting the effect of aggregating data across tasks.
Importantly, these guarantees are established in a model-free, value-based setting with shared representation learning, and therefore complement prior analyses that focus on either single-task offline Q-learning or model-based multitask methods. Our results show that, within this formulation, exploiting shared structure across tasks can improve statistical efficiency without worsening the dominant scaling in the sample size or horizon. Theorems: 1, 2
We further analyzed a downstream offline learning setting in which a novel task shares the representation learned from upstream multitask data. In this case, fixing the learned representation reduces the effective complexity of the hypothesis class for the downstream problem, leading to improved sample efficiency relative to learning both representation and value function from scratch, while retaining the same order of horizon dependence. Overall, our findings demonstrate that Multitask Fitted Q-iterations (MFQI) Borsa et al. (2016) via value-based methods admit controlled finite-sample guarantees under standard assumptions.
@@
MORL (Ishfaq et al. (2024) ) Offline, multitask $\tilde{O}\!\left({H^{2}\lambda}/{\sqrt{nT}}\right)$
REFUEL (Cheng et al. (2022a)) Offline / online, multitask $\tilde{O}\!\left(\varepsilon_{\text{rep}}+{H^{2}\lambda}/{\sqrt{n}}\right)$
Pessimistic Q-learning (Jin et al. (2021)) Offline, single-task $\tilde{O}\!\left({H^{2}\lambda}/{\sqrt{n}}\right)$
@@ Appendix A Proof of Theorem 1a
where $\hat{\phi},\hat{P}^{(1)},\ldots,\hat{P}^{(T)}$ are the output of Algorithm 1 and an appropriete constant $B$. This follows standard proof mechanisms in literature Ishfaq et al. [2024], Agarwal et al. [2020a].
@@ Appendix B Proof of Theorem 1b
We introduce the term $\lambda_{\max}$. Please refer to appendix section F for more detailed analysis of $\lambda_{\max}$. We would now look at the individual terms, in the RHS:
@@ Appendix C Proof of Theorem 1c
Appendix D Proof of theorem-2
@@ Appendix E Note on Raedemacher Complexity
To facilitate a deeper understanding of the bounds in Theorem 2, which leverage Rademacher complexity Bartlett and Mendelson [2002], Shalev-Shwartz and Ben-David [2014] for sharper statistical rates in the downstream task, we provide a brief overview of Rademacher complexity. This measure of function class complexity is a key tool in statistical learning theory, offering tighter generalization bounds than alternatives like VC dimension, especially for non-i.i.d. or structured data settings common in reinforcement learning.
@@ E.2 Generalization Bounds Using Rademacher Complexity
A fundamental result (e.g., from Bartlett and Mendelson, 2002 Bartlett and Mendelson [2002]) states that, with high probability $1-\delta$, for any $g\in\mathcal{G}$ and bounded loss $\ell:[0,1]$,
@@
In summary, Rademacher complexity enables variance-aware, instance-dependent bounds, making it ideal for our sharpened downstream analysis where upstream multitasking effectively lowers the class complexity for the new task. For further reading, see Mohri et al. [2018a] or Shalev-Shwartz and Ben-David [2014].
@@ Appendix F Role of $\lambda_{\max}$ in the Proof
The concentrability coefficient $\lambda_{\max}$ governs how estimation error propagates across Bellman updates in the presence of distribution shift. It appears in Theorems 1 and 2 as a multiplicative factor that quantifies the mismatch between the offline data distribution and the distributions induced by Bellman backups.
Recall that we define
$\lambda_{\max}\triangleq\sup_{h\in[H]}\sup_{\pi}\left|\frac{\mu_{h}^{\pi}}{\mu_{b}}\right|_{\infty},$ (12)
where $\mu_{b}$ is the state–action distribution induced by the behavior policy, and $\mu_{h}^{\pi}$ is the state–action distribution at step $h$ under policy $\pi$. This coefficient upper bounds the worst-case density ratio between the distributions encountered during learning and those present in the offline dataset.
@@
In an ideal on-policy setting, the Bellman optimality operator $T_{h}^{*}$ is a $\gamma$-contraction in an $L_{2}$ norm defined with respect to the same distribution under which errors are measured. However, in the offline setting, the error is evaluated under $\mu_{b}$, while the Bellman operator implicitly involves expectations over next-state distributions induced by the transition kernel and a target policy. This mismatch introduces a change of measure.
Concretely, for any function $f$, we have
$\mathbb{E}_{(s,a)\sim\mu_{h}^{\pi}}\left[f(s,a)^{2}\right]\leq\lambda_{\max}\mathbb{E}_{(s,a)\sim\mu_{b}}\left[f(s,a)^{2}\right],$ (13)
which follows directly from the definition of $\lambda_{\max}$ as a density ratio bound.
Applying this change-of-measure inequality in the Bellman error recursion yields
[…]
Taking square roots, the one-step error propagation satisfies
[…]
Thus, $\sqrt{\lambda_{\max}}$ appears at each step of the recursion, and unrolling this recursion over the horizon leads to polynomial dependence on $\lambda_{\max}$ in the final bounds. For notational simplicity, we sometimes write $\lambda_{\max,\sup}$ to emphasize worst-case concentrability across all steps and policies; this is equivalent to $\lambda_{\max}$ as defined above.
Intuitively, $\lambda_{\max}$ captures how well the behavior policy covers the regions of the state–action space that are important for evaluating or improving the policy. When coverage is good, $\lambda_{\max}$ remains bounded, and the resulting generalization bounds scale polynomially in the horizon and sample size. In contrast, poor coverage leads to large $\lambda_{\max}$, reflecting the intrinsic difficulty of offline reinforcement learning under distribution shift.
This use of concentrability is standard in offline RL theory (e.g., Jin et al. [2021]), and is essential for converting Bellman contraction properties—valid under idealized distributions—into guarantees that hold under the fixed offline data distribution.
Appendix G Experiments under Non-realisability
G.1 Observations under non-realisability
Figure 2: T-Scaling Bound: $O(1/\sqrt{nT})$. As the number of tasks increases (with fixed $n=200$, $H=5$), the average MSE to the true $Q^{*}$ decreases consistently, confirming the multitask efficiency predicted by our bounds.
To empirically validate the theoretical trade-off between statistical efficiency and approximation error, we conduct a scaling study of the MTFQI optimality gap. We simulate $T=5$ tasks in a state space $\mathcal{S}\subset\mathbb{R}^{20}$ where the optimal $Q$-functions are characterized by a low-rank structure of rank $k^{*}=10$. We compare a realizable regime, where our hypothesis class $\mathcal{F}$ matches the environment’s rank, against a misspecified regime where the reward signal is augmented with a high-frequency component $0.8\sin(10\cdot s_{1})$ that lies outside the span of the learned representation $\Phi$. As illustrated in Figure 2, the realizable setting (blue) exhibits a consistent $O(1/n)$ decay in Mean Squared Error, confirming that pooling data across tasks enables the shared encoder to recover the underlying subspace as $n\to\infty$. Conversely, in the non-realizable case (red), the optimality gap initially mirrors the statistical decay but asymptotically plateaus at a constant error floor. This floor represents the inherent Bellman error $\mathcal{E}_{Approx}$, which remains invariant to sample size. This bifurcation numerically confirms our theorem: while the multi-task sample complexity improves the rate of convergence at a $1/\sqrt{nT}$ scale, the final performance is strictly bounded by the fidelity of the representation to the true MDP dynamics. This underscores the necessity of balancing the bottleneck dimension $d$ to mitigate the risk of high-frequency information loss in complex offline environments.